OpenStudy (anonymous):

What is the difference of four and five over eight minus three and one over four in simplest form? 2 1 one and three over eight one and three over four

OpenStudy (anonymous):

@mke

OpenStudy (anonymous):

draw it please

OpenStudy (anonymous):

draw what

OpenStudy (anonymous):

the question, can you?

OpenStudy (anonymous):

@texaschic101 I don't think i can do this one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry, let me know the next one :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there are a whole bunch of people viewing this, i don't know why they're not answering (hint hint people viewing)

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

It is really confusing! I got 1/8 but the next one.... idk

OpenStudy (anonymous):

ok just keep trying plz

OpenStudy (anonymous):

What level is this @keke12147

OpenStudy (anonymous):

wht do u mean?

OpenStudy (anonymous):

are you sure you typed the question right? It's really confusing because of the negative when it says the difference between 4 and 5, thats negative 1 and it messes up the problem

OpenStudy (anonymous):

I think i got it! Give me a moment @keke12147

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It's one and 3 over 8. I'll post the working, don't ask me how because I never came across this type of question! Involved little bit of luck

OpenStudy (anonymous):

What?? is the answer correct?

OpenStudy (anonymous):

idk yet

OpenStudy (anonymous):

Do confirm it! So the working: 4 5/8 - 3 1/4 = 1 and 3/8, since 4 - 3 = 1 and 5/8 - 1/4 = 3/8 i hope u understand :) weird question!